Calculate the acceleration due to gravity on Jupiter. The radius of Jupiter is about 6.99 x 107 m and its mass is 1.9 x 1027 kg.

____m/s^2(downward)
Answer #1

a=Gm/r2

=6.67X10-11 X1.9 x 1027/(6.99 x 107)2

=25.9 m/s2
Answer #2
g = GM / r²

Where g is the acceleration due to gravity, G is the gravitational constant (6.67x10^-11), M is the mass of the object (kg) and r is the radius (m).

g = (6.67*10^-11)*(1.9*10^27) / (6.99*10^7)² = 25.94 m/s²

This is roughly the actual surface acceleration on Jupiter.